# Within the Dynamic Context: Inertia-aware 3D Human Modeling with Pose Sequence

### v1

@@ Abstract
Neural rendering techniques have significantly advanced 3D human body modeling. However, previous approaches often overlook dynamics induced by factors such as motion inertia, leading to challenges in scenarios like abrupt stops after rotation, where the pose remains static while the appearance changes. This limitation arises from reliance on a single pose as conditional input, resulting in ambiguity in mapping one pose to multiple appearances.
In this study, we elucidate that variations in human appearance depend not only on the current frame’s pose condition but also on past pose states. Therefore, we introduce Dyco, a novel method utilizing the delta pose sequence representation for non-rigid deformations and canonical space to effectively model temporal appearance variations. To prevent a decrease in the model’s generalization ability to novel poses, we further propose low-dimensional global context to reduce unnecessary inter-body part dependencies and a quantization operation to mitigate overfitting of the delta pose sequence by the model. To validate the effectiveness of our approach, we collected a novel dataset named I3D-Human, with a focus on capturing temporal changes in clothing appearance under approximate poses. Through extensive experiments on both I3D-Human and existing datasets, our approach demonstrates superior qualitative and quantitative performance. In addition, our inertia-aware 3D human method can unprecedentedly simulate appearance changes caused by inertia at different velocities.
See project page at https://ai4sports.opengvlab.com/Dyco.
@@ 1 Introduction
Our question is whether pose-based LBS processes are sufficient for modeling 3D human motion. To answer this question, let us consider the scenario captured from a static camera view in Fig. 1, wherein a girl spins ($\bm{p}_{1}\to\bm{p}_{3}$) and subsequently stops ($\bm{p}_{3}\to\bm{p}_{4}$). During the cessation of motion, the pose remains essentially unchanged, while the appearance undergoes a transition from strong to weak due to inertia. Clearly, relying solely on the current single pose to model variations in non-rigid human motion and appearance is insufficient, as it overlooks dynamic motion contexts. Consequently, prior works encounter the common occurrence of artifacts stemming from ambiguity in mapping identical pose inputs to different body renderings.
In this work, we propose the Sequence Human method, dubbed Dyco, that significantly advances dynamic human motion modeling by mitigating the ambiguity that arose from misguided reliance on pose conditions. Drawing inspiration from dynamic blend shapes in DMPL [2], we reformulate the non-rigid transformation term to consider past pose sequences, thereby enhancing the model’s ability to generalize to novel pose sequences during testing. Furthermore, we address the oversight of pose-dependent appearance changes in the naive network by introducing a reformulation that incorporates the pose-sequence state, capturing variations in motion-related clothing’s surface appearance under different pose sequences. Note that introducing the concept of pose sequences actually entails a risk, namely, the potential for overfitting due to increased representational capacity. Therefore, we introduce a localized dynamic context encoder to reduce input complexity without sacrificing precision.
In addition to the methodological neglect of dynamic contexts, current datasets fail to accurately capture the associated issues. existing datasets mitigate the issue of appearance variations caused by dynamic contexts by reducing the magnitude of motion (DNA-Rendering Dataset [7]) or wearing tight-fitting garments (ZJU-Mocap Dataset [30], ENeRF-Outdoor Dataset [21] and NHR Dataset [41]). However, real-world human body modeling cannot avoid dealing with large motions and loose clothing. To fully demonstrate the advantages of our inertia-aware 3D human modeling method, we have collected a new dataset of human body data, named I3D-Human. I3D-Human emphasizes significant appearance variations resulting from dynamic contexts, such as skirt swaying and coat tightening, with the aim of facilitating relevant subsequent research.
In summary, our contributions are as follows:
@@
    and I3D-Human Dataset, a self-collected human dataset capturing richer variations of appearance from dynamic contexts. We achieve best results on our I3D-Human Dataset and demonstrate comparable outcomes on other publicly available datasets, compared to state-of-the-art human body modeling works.
@@ 2.3 Modeling Dynamic Context in Human Avatars
The impact of dynamic context on human appearance was discussed in some previous literature. First, there are physical-based approaches which exploited the material properties of human soft tissues and simulated their dynamic simulation using intricate physical models. For example, Larboulette et al. [18] design a mechanical model of a regular limb to simulate the motion of flesh under different accelerations of the bone with the finite element method (FEM). Rohmer et al. [34] underscores the importance of secondary animation for liveliness and presents velocity skinning on top of standard LBS to create dynamic effects such as squash and stretch. However, physical-based approach requires manual design for different materials, which is infeasible to apply to full-body animation. Instead of modeling the underlying physics, another line of work attempts to learn the approximation of dynamics from 4D scans in a data-driven manner. DMPL [2] and Dyna [32] learn to predict soft-tissue deformation from the changing of the body pose. DRAPE [10] incorporates motion history into clothing deformation prediction to animate the draping and flowing effect of clothing. However, these methods are confined to mesh-based representation and deformation and therefore far from high-resolution photo-realistic renderings. On the other hand, current implicit neural avatars are promising for photo-realisim. However, none of them manage to represent the effect of dynamic context and motion inertia on appearance variation. Our work is the first to make the implicit avatar inertia-aware. One reason for the dynamic effect being overlooked is that current datasets, e.g. ZJU-MoCap [30], PeopleSnapshot [3], only focus on tightly-clothed figures with slow and gradual motion. To study the dynamic motion effects that commonly exist in the real scenario, we curate a new dataset of loosely-clothed humans with free and abrupt motion.
@@ 4 Method
4.1 Ambiguities under Identical Pose Conditions
Relying solely on rigid transformations, $T_{R}$, for modeling details of human motion is not sufficient. In many human NeRF works [20, 40], achieving comprehensive motion modeling requires incorporating non-rigid deformations $T_{NR}(\bm{x}_{r},\bm{p})$ conditioned on human poses $\bm{p}$. Some studies [25] introduce additional pose conditions for training the canonical volume to capture variations in per-frame appearance. Nevertheless, our analysis suggests that relying exclusively on pose conditions for non-rigid deformation and canonical volume learning still imposes inherent ambiguities. This is exemplified by varying clothing states of a person under identical poses. In instances where an individual wearing loose clothing abruptly halts after rotation, there is a brief period during which the overall pose remains static, yet inertia induces substantial swinging motion at the specific location $\bm{x}$ on the clothing. Consequently, the non-rigid deformation target $\Delta\bm{x}$ for point $\bm{x}$ becomes ambiguous due to this swinging motion, given their shared condition under identical poses.
Another type of ambiguity arises from the variations in appearance under identical pose conditions, when two observation points corresponding in canonical space exhibit different local normal characteristics. Even under invariant lighting conditions, differences in local normals can impact shading outcomes, consequently affecting the final appearance.
@@ 4.2 Modeling Pose-Sequence Dependent Human Motion
where the sequence length $L$ determines the number of frames in the pose-sequence, while $s$ dictates the level of granularity. Both parameters jointly determine the pose sequence’s window length and motion details of the pose sequence. However, this pose-sequence design is susceptible to overfitting when used as network input, particularly due to its complexity with increasing sequence length $L$.
@@ 4.4 I3D-Human Dataset and Dynamic Human Evaluation
The I3D-Human Dataset: The dataset focuses on capturing variations in clothing appearance under approximately identical poses. Compared with existing benchmarks, we outfit the subjects in loose clothing such as dresses and light jackets and encourage movements involving acceleration or deceleration, such as sudden stops after spinning, swaying, and flapping sleeves. Our capturing equipment consists of 10 DJI Osmo Action cameras, shooting at a frame rate of 100fps while synchronized with an audio signal. The final processed dataset records 10k frames of sequence from 6 subjects in total. More details about the I3D-Human Dataset can be found in our supplemental materials. We will release the dataset for research purposes.
@@ 5.1 Datasets
Given our central objective of imbuing neural implicit avatars with inertia-aware dynamic effects, we mainly examine and compare the methods on the I3D-Human dataset. We conduct experiments on 6 sequences from the dataset, each of which is split into around 600 training frames and 400 test frames. We take 4 camera views for training and the rest for testing. In addition to modeling dynamic motion, it remains crucial to maintain the performance of the model in the prototypical scenario of tightly-clothed subjects in slow and controlled motion. Therefore, we also compare our method with prior work on the most commonly used ZJU-MoCap [31] benchmark. ZJU-MoCap comprises 9 multi-view RGB sequences of human motion. Following the training setup in previous work [31, 27, 29], we use 4 cameras for training and the rest 19 cameras for testing. Each sequence is divided into two consecutive clips, one for training and the other for testing unseen poses.
5.2 Evaluation
We compare the methods on two test splits, namely novel view and novel pose. Novel-view test set evaluate the observed poses in the training frames from novel camera views, assessing whether the model can learn a physically realistic 3D human under the training poses from the 2D RGB observations. Novel-pose test set renders unobserved poses from novel camera views, examining whether the learned 3D avatars can be plausibly animated under unseen poses. For evaluation metrics, we adopt traditional per-frame metrics including PSNR, SSIM, and LPIPS [43]. As explained in Section Sec. 4.4, we also propose a novel motion-based metric termed DME to measure the accuracy of motion dynamics in the I3D-Human dataset.
@@ 5.3 Comparison with State of the Arts
We compare with state-of-the-art human avatars, including Neural Body [28], AniNeRF [29], AniSDF [29], HumanNeRF [40] and 3DGS-Avatar [33]. The last two models are initially proposed for the monocular training setting so we adapt them to our multi-view setting. It is worth noting that the concurrent 3DGS-Avatar harnesses the advanced capabilities of Gaussian splatting representation [16].
I3D-Human: We show the comparisons on I3D-Human in Tabs. 1, 3 and 4. Our approach not only surpasses state-of-the-arts in terms of standard frame-wise similarity metrics but also achieves significantly lower DME, indicating its effectiveness in modeling inertia-induced dynamics. This improvement remains consistent across both novel view and novel pose test sets, indicating that the learned correlation between dynamic context and appearance variation in training frames can be extended to unseen poses. We will delve deeper into the generalization later.
Figure 3: Qualitative comparison on the novel view of I3D-Human dataset. Our method can render distinct appearance and deformation of the drape and wrinkles across various motion contexts. In contract, HumanNeRF [40] and 3DGS-Avatar [33] struggle to capture the precise details. The other three baselines [29, 28] exhibit noticeable artifacts.
@@
ZJU-MoCap: To show Dyco maintains the capability of modeling motion with minimal inertia influence, we compare it with state-of-the-arts on ZJU-MoCap and report the metrics in Tab. 2. Despite the deliberate minimization of inertia effects in human motion within ZJU-MoCap, our method achieves the best LPIPS scores in the novel-view renderings of the training poses. We find that even for ZJU-MoCap, in some training frames with similar static poses, variations in cloth wrinkles can still occur due to divergent past motion trajectories. We visualize this in the supplementary. In the novel pose test, our method only slightly underperforms HumanNeRF and 3DGS-Avatar. Qualitative comparison can be found in the supplementary.
Figure 5: The ZJU-MoCap dataset also exhibits ambiguous mapping from the static pose to various appearances. At two frames when the subject has similar poses, the clothing wrinkles can differ due to distinct past motion. Our method can reflect the variation, while HumanNeRF [40] generates similar patterns. The results are novel view renderings.
@@ 5.4 Ablations
PSNR SSIM $\textrm{LPIPS}^{*}$ DME PSNR SSIM $\textrm{LPIPS}^{*}$ DME
@@
PSNR SSIM $\textrm{LPIPS}^{*}$ DME PSNR SSIM $\textrm{LPIPS}^{*}$ DME
@@
We conduct ablation experiments to validate the effectiveness of our designed components. In Tab. 3, we investigate the impact of pose condition $\bm{p}_{i}$, delta pose condition $S_{\Delta{\bm{p}}}$, and the localized spatial dependency in dynamic context encoder.
Pose condition $\bm{p}_{i}$ provides global information of the current frame for network training, whose removal is accompanied by a foreseeable decrease in performance.
Delta pose condition $S_{\Delta{\bm{p}}}$ incorporates temporal human inertia knowledge, which we believe is the core design for resolving ambiguities related to dynamic context. Discarding this component prevents the network from leveraging inertia information and degrades the performance.
Localization in dynamic context encoder reduces the complexity of $S_{\Delta{\bm{p}}}$ input by masking redundant dependencies. Without this module, the overfitted network will struggle in generalizing to novel poses.
Conditional input to $T_{\textrm{NR}}$ and $F_{c}$: In Sec. 4.1, we analyze two types of dynamic context related ambiguities and apply delta pose $S_{\Delta{\bm{p}}}$ to both non-rigid transformation and canonical volume. In Tab. 4, we further demonstrate the existence of such ambiguities through stepwise ablation of the input $S_{\Delta{\bm{p}}}$ in these two modules.
@@ 5.5 Generalization to Novel Dynamic Context
Varying velocity: We scale $S_{\Delta{\bm{p}}}$ with a factor $\alpha$ ranging from 00 to $2$, where $\alpha>1$ speeds up motion. The scaled $\alpha{S_{\Delta{\bm{p}}}}$ is used as the sequence condition. Fig. 6 demonstrates Dyco’s ability to accurately portray varying drape lifting amplitudes with different spinning speeds, with larger $\alpha$ yielding greater amplitudes.
Varying acceleration: We simulate an abrupt stop effect by interrupting poses midway and making subsequent poses identical. The modified pose sequence is then used as input condition for the trained model. As depicted in Fig. 7, Dyco, when conditioned on the novel pose sequence, accurately animates the falling of the hem. In contrast, the model without pose sequence inputs fails to depict this motion.

### v2

@@ Abstract
Neural rendering techniques have significantly advanced 3D human body modeling. However, previous approaches overlook dynamics induced by factors such as motion inertia, leading to challenges in scenarios where the pose remains static while the appearance changes, such as abrupt stops after spinning. This limitation arises from conditioning on a single pose, which leads to ambiguity in mapping one pose to multiple appearances.
In this study, we elucidate that variations in human appearance depend not only on the current frame’s pose condition but also on past pose states. We introduce Dyco, a novel method that utilizes the delta pose sequence to effectively model temporal appearance variations. To mitigate overfitting to the delta pose sequence, we further propose a localized dynamic context encoder to reduce unnecessary inter-body part dependencies. To validate the effectiveness of our approach, we collect a novel dataset named I3D-Human, focused on capturing temporal changes in clothing appearance under similar poses. Dyco significantly outperforms baselines on I3D-Human and achieves comparable results on ZJU-MoCap. Furthermore, our inertia-aware 3D human method can unprecedentedly simulate appearance changes caused by inertia at different velocities. The code, data and model are available at our project website at https://ai4sports.opengvlab.com/Dyco.
@@ 1 Introduction
Our question is whether pose-based LBS processes are sufficient for modeling 3D human motion. To answer this question, let us consider the scenario captured from a static camera view in Fig. 1, wherein a girl spins ($\bm{p}_{1}\to\bm{p}_{3}$) and then stops ($\bm{p}_{3}\to\bm{p}_{4}$). As she stops, her pose stays mostly the same, but the drape of her dress noticeably falls. Therefore, solely relying on the current single pose is insufficient to model variations in non-rigid human motion and appearance, as it ignores dynamic motion contexts. This limitation causes previous methods to struggle with mapping identical pose inputs to different appearance, resulting in artifacts in the scenario described above.
In this work, we propose a neural avatar model that accounts for dynamic context, dubbed Dyco, which significantly advances dynamic human motion modeling by resolving the ambiguity that arises from relying solely on static poses. Drawing inspiration from dynamic blend shapes in DMPL [2], we encode pose sequences as condition in addition to the current static pose. We compute the changes, or deltas, in pose parameters over the pose trajectory and input the delta sequence into the non-rigid deformation field and canonical volume. Since the increased input dimension poses a risk of overfitting, we design a localized dynamic context encoder to reduce input complexity without sacrificing information. The dynamic context encoder eliminates spurious dependencies between different body parts and efficiently aggregates spatial and temporal information.
In addition to the methodological neglect of dynamic contexts, current datasets fail to accurately capture the associated issues. existing datasets mitigate the issue of appearance variations caused by dynamic contexts by reducing the magnitude of motion (DNA-Rendering Dataset [7]) or wearing tight-fitting garments (ZJU-Mocap Dataset [30], ENeRF-Outdoor Dataset [21] and NHR Dataset [41]). However, real-world human body modeling cannot avoid dealing with large motions and loose clothing. To fully demonstrate the advantages of our inertia-aware 3D human modeling method, we have collected a new dataset of human body data, named I3D-Human. I3D-Human emphasizes significant appearance variations resulting from dynamic contexts, such as skirt swaying and hem dropping, with the aim of facilitating relevant future research. In summary, our contributions are as follows:
@@
    I3D-Human Dataset, a self-collected human dataset capturing richer variations of appearance from dynamic contexts. We achieve best results on our I3D-Human Dataset and demonstrate comparable performance on ZJU-Mocap, compared to state-of-the-art human body modeling methods.
@@ 2.3 Modeling Dynamic Context in Human Avatars
The impact of dynamic context on human appearance was discussed in some previous literature. First, there are physical-based approaches which exploited the material properties of human soft tissues and simulated their dynamic simulation using intricate physical models. For example, Larboulette et al. [18] design a mechanical model of a regular limb to simulate the motion of flesh under different accelerations of the bone with the finite element method (FEM). Rohmer et al. [34] underscores the importance of secondary animation for liveliness and presents velocity skinning on top of standard LBS to create dynamic effects such as squash and stretch. However, physical-based approach requires manual design for different materials, which is infeasible to apply to full-body animation. Instead of modeling the underlying physics, another line of work attempts to learn the approximation of dynamics from 4D scans in a data-driven manner. DMPL [2] and Dyna [32] learn to predict soft-tissue deformation from the changing of the body pose. However they can only model naked human bodies. DRAPE [10] incorporates motion history into clothing deformation prediction to animate the draping and flowing effect of clothing. However, it requires the clothing template mesh and its proficiency is highly limited to the types of clothing it is trained on. Also, these methods are confined to mesh-based representation and deformation and thus far from high-resolution photo-realistic renderings.
On the other hand, current implicit neural avatars are promising for photo-realisim. However, none of them manage to represent the effect of dynamic context and motion inertia on appearance variation. Our work is the first to make the implicit avatar inertia-aware. One reason for the dynamic effect being overlooked is that current datasets, e.g. ZJU-MoCap [30], PeopleSnapshot [3], only focus on tightly-clothed figures with slow and gradual motion. To study the dynamic motion effects that commonly exist in the real scenario, we curate a new dataset of loosely-clothed humans with free and abrupt motion.
@@ 4 Method
4.1 Ambiguities under Identical Pose Conditions
Relying solely on rigid transformations, $T_{R}$, for modeling details of human motion is not sufficient. In many human NeRF works [20, 40], achieving comprehensive motion modeling requires incorporating non-rigid deformations $T_{NR}(\bm{x}_{r},\bm{p})$ conditioned on human poses $\bm{p}$. Some studies [25] introduce additional pose conditions for training the canonical volume to capture variations in per-frame appearance. Nevertheless, our analysis suggests that relying exclusively on pose conditions for non-rigid deformation and canonical volume learning still imposes inherent ambiguities. This is exemplified by varying clothing states of a person under identical poses. In instances where an individual wearing loose clothing abruptly halts after rotation, there is a brief period during which the overall pose remains static, yet inertia induces substantial swinging motion at the specific location $\bm{x}$ on the clothing. Consequently, the non-rigid deformation target $\Delta\bm{x}$ for point $\bm{x}$ becomes ambiguous due to this swinging motion, given their shared condition under identical poses. Another type of ambiguity arises from the variations in appearance under identical pose conditions, when two observation points corresponding in canonical space exhibit different local normal characteristics. Even under invariant lighting conditions, differences in local normals can impact shading outcomes, consequently affecting the final appearance.
@@ 4.2 Modeling Pose-Sequence Dependent Human Motion
where the sequence length $L$ is the number of frames in the pose-sequence and $s$ is the level of granularity. Both parameters jointly determine the pose sequence’s window length and motion details of the pose sequence. However, this pose-sequence design is susceptible to overfitting when used as network input, particularly due to its complexity with increasing sequence length $L$.
@@ 4.4 I3D-Human Dataset and Dynamic Human Evaluation
The I3D-Human Dataset: The dataset focuses on capturing variations in clothing appearance under approximately identical poses. Compared with existing benchmarks, we outfit the subjects in loose clothing such as dresses and light jackets and encourage movements involving acceleration or deceleration, such as sudden stops after spinning, swaying, and flapping sleeves. Our capturing equipment consists of 10 DJI Osmo Action cameras, shooting at a frame rate of 100fps while synchronized with an audio signal. The final processed dataset records 10k frames of sequence from 6 subjects in total. More details about the I3D-Human Dataset can be found in our supplemental materials.
@@ 5.1 Datasets
Given our central objective of imbuing neural implicit avatars with inertia-aware dynamic effects, we mainly examine and compare the methods on the I3D-Human dataset. We conduct experiments on 4 sequences from the dataset and split each sequence into around 600 training frames and 400 test frames. We take 4 cameras for training and the rest for testing. As it remains crucial to maintain performance in the scenario of tightly-clothed subjects in slow and controlled motion, we also compare our method with others on the mostly used ZJU-MoCap [31] benchmark. ZJU-MoCap comprises 9 multi-view RGB sequences. Following the training setup in previous work [31, 27, 29], we use 4 cameras for training and the rest 19 cameras for testing. Each sequence is divided into two consecutive clips, one for training and the other for testing unseen poses.
@@
5.2 Evaluation
We compare the methods on two test splits, namely novel view and novel pose. Novel-view test set evaluates the observed poses in the training frames from novel camera views, assessing whether the model can learn a physically realistic 3D human under the training poses from the 2D RGB observations. Novel-pose test set renders unobserved poses from novel camera views, examining whether the learned 3D avatars can be plausibly animated under unseen poses. For evaluation metrics, we adopt traditional per-frame metrics including PSNR, SSIM, and LPIPS [43]. As explained in Sec. 4.4, we also propose a novel motion-based metric termed DME to measure the accuracy of motion dynamics rendering.
@@ 5.3 Comparison with State of the Arts
We compare with state-of-the-art counterparts, including Neural Body [28], AniNeRF [29], AniSDF [29], HumanNeRF [40] and 3DGS-Avatar [33]. The last two models are initially proposed for the monocular training setting so we adapt them to our multi-view setting. Note that the concurrent 3DGS-Avatar harnesses the advanced capabilities of Gaussian splatting representation [16].
I3D-Human: We show the comparisons on I3D-Human in Tabs. 1, 3 and 4. Our approach not only surpasses state-of-the-arts in terms of standard frame-wise similarity metrics but also achieves significantly lower DME, indicating its effectiveness in modeling inertia-induced dynamics. This improvement remains consistent across both novel view and novel pose test sets, indicating that the learned correlation between dynamic context and appearance variation in training frames can be extended to unseen poses. We will delve deeper into the generalization performance later.
Figure 3: Qualitative comparison on the novel view of I3D-Human dataset. Our method can render distinct appearance and deformation across various motion contexts. In contrast, HumanNeRF [40] and 3DGS-Avatar [33] struggle to capture the precise details. The other three baselines [29, 28] exhibit noticeable artifacts.
@@
ZJU-MoCap: To show Dyco maintains the capability of modeling motion with minimal inertia influence, we compare it with state-of-the-arts on ZJU-MoCap and report the metrics in Tab. 2. Despite the deliberate minimization of inertia effects in human motion within ZJU-MoCap, our method achieves the best LPIPS scores in the novel-view test set. Even for ZJU-MoCap, in some training frames with similar static poses, variations in cloth wrinkles can still occur due to divergent past motion trajectories. We show that Dyco outperforms other methods in capturing the variation in Fig. 5. In the novel pose test, our method only slightly underperforms HumanNeRF and 3DGS-Avatar. Qualitative comparison can be found in the supplementary material.
Figure 5: The ZJU-MoCap dataset also exhibits ambiguous mapping from the static pose to various appearances. At two frames when the subject has similar poses, the clothing wrinkles can differ due to distinct past motion. Our method can reflect the variation, while HumanNeRF [40] generates similar patterns.
@@ 5.4 Ablations
[…]
@@
[…]
@@
Pose condition $\bm{p}_{i}$ provides global spatial information for the current frame during training, and its removal results in decreased performance, see  Tab. 3.
Delta pose condition $S_{\Delta{\bm{p}}}$ incorporates temporal human inertia knowledge, which we believe is the core design for resolving ambiguities related to dynamic context. Discarding this component prevents the network from leveraging inertia information and degrades the performance, see Tab. 3.
Localization in dynamic context encoder reduces the complexity of $S_{\Delta{\bm{p}}}$ input by masking redundant dependencies. Without this module, the overfitted network will struggle in generalizing to novel poses, see Tab. 3.
Conditional input to $T_{\textrm{NR}}$ and $F_{c}$: In Sec. 4.1, we analyze two types of dynamic context related ambiguities and propose to input delta pose $S_{\Delta{\bm{p}}}$ to both non-rigid transformation and canonical volume. Tab. 4 validates that inputting $S_{\Delta{\bm{p}}}$ into both modules helps resolve the ambiguities.
@@ 5.5 Generalization to Novel Dynamic Context
Varying velocity: We scale $S_{\Delta{\bm{p}}}$ with a factor $\alpha$ ranging from 00 to $2$, where $\alpha>1$ speeds up motion and use $\alpha{S_{\Delta{\bm{p}}}}$ as the sequence condition. Fig. 6 demonstrates Dyco’s ability to accurately portray varying drape lifting amplitudes with different spinning speeds, with larger $\alpha$ yielding greater amplitudes.
Varying acceleration: We simulate an abrupt stop effect by interrupting poses midway and making subsequent poses identical. The modified pose sequence is then used as the input condition. As shown in Fig. 7, Dyco, when conditioned on the novel pose sequence, accurately animates the falling of the hem. In contrast, the model without pose sequence inputs fails to depict this motion.